OpenStudy (anonymous):

A circle is inscribed in a square. The diameter of the circle is 12.4 mm. Find the area of the region that is outside of the circle and inside the square. Round the answer to the nearest tenth.

OpenStudy (anonymous):

Well, you have to find the area of the circle Then find the area of the square Then take the area of the square away from the area of the circle

OpenStudy (anonymous):

radius of circle = 1/2 D = 6.2mm

OpenStudy (anonymous):

area of circle = pi * r^2 = 120.76 mm